A 220V, 50 Hz, 6 pole, Y connected alternator with resistance 0.06 Ω per phase gave

the following data for open circuit and Short circuit characteristics.


1
Expert's answer
2021-07-15T02:48:45-0400

The induced e.m.f per phase can be calculated as follows

Armature current

"I_a= \frac{S}{\sqrt{3}*V_L}= \frac{10*10^6}{\sqrt{3}*220} = 26243.19 A\\"

Synchronous impedance,

"Z_s=R_a+jX_s = 0.07+j0.06 \Omega = 1.901\angle 87.89^0 \Omega"